Stallone’s Italian Restaurant announced an ROA of 9.80% and profit margin ratio of 19.50% in 2012. This company had a total of $20 million total assets at the end of 2012. Based on the information given, calculate the total revenue generated in 2012 for Stallone’s Italian Restaurant.
Answer #1

Let revenue be x

ROA = Net income/Total assets

(x * 0.195)/20,000,000 = 0.098

x * 0.195 = 1,960,000

x = 1,960,000/0.195

= 10,051,282